if |z|=1 find the value of arg(z)-arg(z+1)-arg(z-1)

if |z|=1 find  the value of  arg(z)-arg(z+1)-arg(z-1)

Grade:11

1 Answers

Deepak Kumar Shringi
6 years ago
LET Z= i and now solve it you can put values to solve questions fast
arg(i)=pie/2
arg (1+i)=pie/4
arg (i-1)=3pie/4 so answer will be= -pie/2